OpenStudy (anonymous):

A cube is cut into smaller cubes such that each edge if cube is increased by 50% ,find the % increase in surface area?

OpenStudy (anonymous):

(A) 200% (B) 100% (C) 125% options

OpenStudy (marigirl):

each edge OF cube is increased by 50%?

OpenStudy (anonymous):

yup dr

OpenStudy (anonymous):

well, then increase of 50% means that new edge is 1.5 times the older one.

OpenStudy (anonymous):

how u get that @decidedlypretentious

OpenStudy (anonymous):

\[x + 50/100 x = 150/100x\]

OpenStudy (anonymous):

btw, have a lookie at your question again, does it say decreased or increased? cause if its cut into smaller cubes... ehh

OpenStudy (anonymous):

question is lyk dat only

OpenStudy (anonymous):

Assuming the question doesn't mean the first bit.

OpenStudy (anonymous):

I'd say then calculate surface area for both cases. (Final SA-Initial SA)/Initial SA * 100 should give you the answer

OpenStudy (anonymous):

okayyyy.....do you ne shortcut key for this typ of questions.

OpenStudy (anonymous):

n yeah definately it will give the ans..thnkyou so mch :)

OpenStudy (marigirl):

Let the edge = a cm So increase by 50 % = a + a/ 2 = 3a / 2 TSA of original cube = 6 a 2 TSA of new cube = 6 ( 3a/ 2) 2 = 6 x 9a 2 / 4 = 54a 2 / 4 = 13.5a 2 Increase in area = 13.5a 2 - 6 a 2 = 7.5a 2 Increase % = ( 7.5a 2 / 6 a 2 ) x 100 = 125%

OpenStudy (anonymous):

u r genius...thnkyou so mch @marigirl